William is 6yrs older than robert. in two years he will be twice as old as robert. find their ages.can you help me with this problem? thanks
Answer by stanbon(75887)

William is 6yrs older than robert. in two years he will be twice as old as robert. find their ages.can you help me with this problem? thanks
--------------------
NOW Data:
Robert is x yrs. old; William is x+6 yrs old.
2 YRS. from NOW:
Robert will be x+2 yrs. old; William will be x+8 yrs. old
EQUATION:
In 2 years William will be twice as old as Robert.
x+8=2(x+2)
x+8=2x+4
x=4 (Robert's age NOW
x+6=10 (William's age NOW)
Cheers,
Stan H.